OpenStudy (anonymous):

a particlestarting from rest moves with constant acceleration. if it takes 5.0 s to reach the speed 180. km/h find the average velocity during this period

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (akashdeepdeb):

\[Average Velocity = \frac{ v - u }{ \Delta t }\] You have Δt = 5s u = 0m/s [Starts from rest] v = 50m/s [Convert it into SI units] Plud the values in and you'll get the Average Velocity. Understood? :)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (goformit100):

Madam @pinki , @AkashdeepDeb gave the correct answer.

OpenStudy (theeric):

I agree with @AkashdeepDeb , also! :) So, \(v\) is the magnitude of the velocity. There is a symbol commonly used to denote the average of something. \(\bar v\) is the average velocity. So, if you want to use that in your equations, or \(v_\text{ave}\), you'll have formal variables! For coding the equations, \(\bar v\) = `\(\bar v\)` \(v_\text{ave}\) = `\(v_\text{ave}\)` Just thought I'd throw some extra information out there!